Six grapefruit cost as much as ten oranges. The cost of a dozen grapefruit and two dozen oranges is $33. How many cents does one orange cost?
Answer by swincher4391(1107)

Let g be the cost of grapefruit. Let o be the cost of oranges.
6g = 10o
12g + 2o = 33
6g -10o = 0
12g +2o = 33
Multiply the 1st equation by -2.
-12g +20o = 0
12g + 2o = 33
Add the two together:
0g + 22o = 33
o = 3/2 = $1.50
6g = 10(1.50)
6g = 15
g = 15/6 = 5/2 = $2.50
Oranges cost $1.50
Grapefruit costs $2.50
Hope this helped!